# Local Topological Constraints on Berry Curvature in Spin–Orbit Coupled BECs

### v2

@@ Abstract
We establish a local topological obstruction to flattening Berry curvature in spin-orbit-coupled Bose-Einstein condensates (SOC BECs), valid even when the global Chern number vanishes. For a generic two-component SOC BEC, the extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ carries a natural metric connection $\nabla^{C}$ whose torsion 3-form encodes the synthetic gauge fields. Its harmonic part defines a mixed cohomology class
@@ 1 Introduction
The first main task of the paper is to construct, in this space, a natural geometric structure to which the theorem of [12] applies. We introduce a metric $g_{M}$ of Kaluza–Klein type, which couples the base metric on $T^{2}_{\mathrm{BZ}}$ to the $U(1)$-valued connection forms that encode the synthetic gauge fields of the SOC BEC. On $(M,g_{M})$ we define a metric connection $\nabla^{C}$ whose torsion three-form is precisely the physical three-form
@@ 2 Geometric Construction of the Extended Parameter Space
$M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}.$
@@ 2.2 Metric connection with prescribed torsion
Define the three‑form
$T^{b}:=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)}\in\Omega^{3}(M),$ (2.2)
where $F^{(\pm)}\wedge A^{(\pm)}$ are pulled back to $M$ via $\pi$ (we keep the same notation for the pull‑backs). $T^{b}$ is the trace $\operatorname{Tr}(\mathcal{F}\wedge\mathcal{A})$ with respect to the Killing form on $\mathfrak{u}(1)_{+}\oplus\mathfrak{u}(1)_{-}$, where $\mathcal{A}=A^{(+)}\otimes t_{+}+A^{(-)}\otimes t_{-}$ and $\mathcal{F}=F^{(+)}\otimes t_{+}+F^{(-)}\otimes t_{-}$.
@@ 2.3 Parallel forms in the Kaluza–Klein geometry
Lemma 2.4 (Non‑parallelism of $d\phi_{\pm}$).
Assume $A^{(\pm)}$ are not identically zero. Then the one‑forms $d\phi_{+},d\phi_{-}$ are not parallel with respect to $g_{M}$.
@@
We compute the covariant derivative $\nabla^{LC}d\phi_{+}$ in local coordinates $(k_{x},k_{y},\phi_{+},\phi_{-})$. The non‑zero Christoffel symbols involving the index $\phi_{+}$ are
$\Gamma^{\phi_{+}}_{k_{x}\phi_{+}}=\frac{1}{2}g^{\phi_{+}\rho}\bigl(\partial_{k_{x}}g_{\phi_{+}\rho}+\partial_{\phi_{+}}g_{k_{x}\rho}-\partial_{\rho}g_{k_{x}\phi_{+}}\bigr).$
Since $g_{k_{x}\phi_{+}}=A^{(+)}_{k_{x}}$ is independent of $\phi_{+}$, we have $\partial_{\phi_{+}}g_{k_{x}\phi_{+}}=0$. Moreover $g_{\phi_{+}\phi_{+}}=1+(A^{(+)}_{k_{x}})^{2}+(A^{(+)}_{k_{y}})^{2}$ and $g_{\phi_{+}\phi_{-}}=A^{(+)}_{k_{x}}A^{(-)}_{k_{x}}+A^{(+)}_{k_{y}}A^{(-)}_{k_{y}}$. Consequently,
[…]
which is non‑zero for generic non‑constant potentials $A^{(\pm)}$. Hence $(\nabla^{LC}_{\partial_{k_{x}}}d\phi_{+})(\partial_{\phi_{+}})=-\Gamma^{\phi_{+}}_{k_{x}\phi_{+}}\neq 0$, showing that $d\phi_{+}$ is not parallel. The same argument applies to $d\phi_{-}$. ∎
Corollary 2.5 (Trivial parallel‑form spaces on the fibres).
For a generic SOC BEC with non‑trivial Berry curvature $F^{(\pm)}\neq 0$,
$\mathcal{P}_{1}\bigl(S^{1}_{\phi_{\pm}},g_{M}|_{\mathrm{fibre}}\bigr)=\{0\}.$
More precisely, any fibre‑tangent 1‑form that is parallel with respect to the full metric $g_{M}$ must be of the form $\xi=c_{+}d\phi_{+}+c_{-}d\phi_{-}$ with constants $c_{+},c_{-}$. However, Lemma 2.4 shows that for non‑constant potentials $A^{(\pm)}$, $\nabla^{LC}d\phi_{\pm}$ has non‑vanishing components along the base directions; consequently no non‑zero linear combination can be parallel on the whole manifold $M$. If $A^{(\pm)}$ are constant but $F^{(\pm)}\neq 0$, they are not closed, hence still not harmonic and therefore not parallel (on a compact Riemannian manifold every parallel form is harmonic).
@@ 3 Cohomology of the Extended Space and the Mixed Tensor Rank
With the geometric structure $(M,g_{M},\nabla^{C})$ established in Section 2, we now analyze the cohomological data that enter the general lower bound. Recall that the torsion three-form $T^{\flat}\in\Omega^{3}(M)$ is not necessarily closed, but its de Rham cohomology class $T^{\flat}]\in H^{3}(M;\mathbb{R})$ is well defined. Let $\omega_{h}$ be the unique $g_{M}$-harmonic representative of this class; then
@@ 3.1 Künneth decomposition of $H^{3}(M)$
For the harmonic representative $\omega_{h}$ of our torsion class, the Hodge decomposition with respect to $g_{M}$ yields
$\omega_{h}=\alpha_{+}\wedge d\phi_{+}+\alpha_{-}\wedge d\phi_{-}+d\eta,$
where $\alpha_{\pm}\in\mathcal{H}^{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})$ are $g_{\mathrm{BZ}}$-harmonic two-forms on the Brillouin torus, and $\eta\in\Omega^{2}(M)$ is a coexact two-form. The term $d\eta$ is exact and therefore orthogonal to harmonic forms in the $L^{2}$-inner product; it does not contribute to the cohomology class $\omega]=[\omega_{h}]$.
@@
The representatives $d\phi_{\pm}$ are harmonic with respect to the standard metric on the circles, but they are not the harmonic representatives of the fundamental classes with respect to the full Kaluza–Klein metric $g_{M}$; the $g_{M}$-harmonic representatives are the horizontal forms $\Theta^{(\pm)}=d\phi_{\pm}+A^{(\pm)}$ introduced in (2.1). Because $d\phi_{\pm}]=[\Theta^{(\pm)}]$ in $H^{1}(S^{1}_{\phi_{\pm}})$, the cohomological decomposition (3.1) is unchanged.
@@ 3.2 Mixed tensor rank of the torsion class
For the class $\omega]$ coming from the physical torsion, the two-forms $\alpha_{+},\alpha_{-}$ in (3.1) are precisely the (cohomology classes of the) Berry curvature two-forms $F^{(+)},F^{(-)}$ of the two independent $U(1)$ sectors. A key observation is that $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$. Therefore, any two non-zero classes in this space are necessarily cohomologically proportional. In a generic two-component SOC BEC, both Berry curvature classes $F^{(+)}]$ and $F^{(-)}]$ are non-zero; consequently, there exists a non-zero constant $\lambda\in\mathbb{R}$ such that
$F^{(-)}]=\lambda[F^{(+)}]\quad\text{in }H^{2}(T^{2}_{\mathrm{BZ}}).$
@@ 3.3 Vanishing of the obstruction kernel $\mathcal{K}$
In our geometry, we have $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\,\mathrm{vol}_{\mathrm{BZ}}$ (the space of parallel two-forms on the flat torus). The crucial observation, already made in Section 2.3, is that for the Kaluza–Klein metric $g_{M}$ the spaces of parallel one-forms on the phase circles are trivial:
$\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},\,g_{M}|_{\text{\rm fibre}})=\{0\},$
because any $g_{M}$-parallel one-form tangent to a fibre would have to be a constant multiple of $\Theta^{(\pm)}=d\phi_{\pm}+A^{(\pm)}$, and $\Theta^{(\pm)}$ is not closed (hence not of the form $d\phi_{\pm}$) unless $F^{(\pm)}=0$. Since the Berry curvatures $F^{(\pm)}$ are generically non-zero, we indeed have $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}})=\{0\}$ in the physical geometry.
By Lemma 4.5 (with $A^{(\pm)}$ non‑flat), the space of parallel 1‑forms on the phase circles vanishes: $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{M})=\{0\}$. Consequently the tensor products $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}})$ and $\mathcal{P}_{1}(S^{1}_{\phi_{+}})\otimes\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})$ (the latter appears in $\ker\widetilde{\Psi}_{p}$) are zero. Thus

### v3

@@ Abstract
We establish a local topological obstruction to flattening Berry curvature in spin-orbit-coupled Bose-Einstein condensates (SOC BECs), valid even when the global Chern number vanishes. For a generic two-component SOC BEC, the extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ carries a natural metric connection $\nabla^{C}$ whose torsion 3-form encodes the synthetic gauge fields. Under the physically relevant assumption of constant Berry curvatures, the harmonic part of this torsion defines a mixed cohomology class
@@ 1 Introduction
Clarification of methodological approach.
While Theorem 5.2 of [12] is formally stated for genuine Riemannian product metrics $g=g_{1}\oplus g_{2}$, the physical metric $g_{M}$ in our SOC BEC model is a Kaluza–Klein metric that is not globally a Riemannian product when the Berry curvatures are non-zero. We do not apply the theorem as a black box. Instead, we adapt its proof to our geometry by: (i) using the pointwise orthogonal splitting $\mathcal{H}_{p}\oplus\mathcal{V}_{p}$ to define obstruction maps $\Psi_{p},\widetilde{\Psi}_{p}$ analogously to those in [12]; (ii) showing that the harmonic representative of the torsion class decomposes appropriately; and (iii) proving that the kernel computation leading to $\mathcal{K}=0$ remains valid. This adaptation, supplemented by a deformation argument ($g_{\varepsilon}\to g_{M}$) and lower-semicontinuity, rigorously yields the bound for the physical metric $g_{M}$.
@@
The first main task of the paper is to construct, in this space, a natural geometric structure to which the theorem of [12] applies. We introduce a metric $g_{M}$ of Kaluza–Klein type, which couples the base metric on $T^{2}_{\mathrm{BZ}}$ to the $U(1)$-valued connection forms that encode the synthetic gauge fields of the SOC BEC. For simplicity and to make the geometric arguments fully explicit, we will work in the experimentally relevant setting where the Berry curvatures are constant multiples of the Brillouin–zone volume form. This covers a wide range of SOC BEC realizations and allows us to write the harmonic representative of the torsion class in closed form. On $(M,g_{M})$ we define a metric connection $\nabla^{C}$ whose torsion three-form is precisely the physical three-form
@@ 2 Geometric Construction of the Extended Parameter Space
$M=T_{\mathrm{BZ}}^{2}\times S_{\phi_{+}}^{1}\times S_{\phi_{-}}^{1}.$
Global structure of the extended parameter space.
The extended parameter space is the total space $M$ of a principal $U(1)_{+}\times U(1)_{-}$-bundle over the Brillouin torus, equipped with the synthetic connections $A^{(+)},A^{(-)}$. As a smooth manifold, $M$ is diffeomorphic to the four-torus $T^{4}$. This diffeomorphism allows us to identify $M$ with a product $T^{2}_{\rm BZ}\times T^{2}_{\rm fibre}$ for the purpose of applying topological results, where $T^{2}_{\rm fibre}\cong S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$.
The bundle is trivial if and only if each individual Chern class vanishes: $c^{(+)}_{1}=0$ and $c^{(-)}_{1}=0$. The physically interesting regime of compensated topology, where $c^{(+)}_{1}=-c^{(-)}_{1}\neq 0$, corresponds to a non-trivial bundle. However, the central results of this paper—constructed from the globally defined metric $g_{M}$ and connection forms $\Theta^{(\pm)}$—do not require bundle triviality, only the global geometric structures on $M\cong T^{4}$. The condition $c^{\rm tot}_{1}=0$ ensures the existence of a global section for the associated $S^{1}$-bundle related to the diagonal subgroup, which is sufficient for the physical interpretation of separable phase variables in a global gauge.
This is precisely the regime where conventional global invariants (like the total Chern number) fail to detect local geometric obstructions, and where the cohomological lower bound developed below becomes particularly relevant.
If the total Chern number were non-zero, the parameter space would be a non-trivial fibre bundle, and the analysis would require working in local trivializations; the topological lower bound would then hold locally in each gauge patch. Crucially, even when the bundle is non-trivial ($c^{(\pm)}_{1}\neq 0$), the total space $M$ remains diffeomorphic to $T^{4}$, and all geometric structures—the metric $g_{M}$, the connection forms $\Theta^{(\pm)}$, and the orthogonal splitting $\mathcal{H}\oplus\mathcal{V}$—are globally defined. Therefore, our geometric analysis applies to both trivial and non-trivial bundles, provided the total Chern number vanishes. This precisely captures the physically important situation of compensated topological sectors.
Mathematical framework.
The geometric construction that follows requires only that:
  1. 1.
    The total space $M$ is diffeomorphic to $T^{4}$ (true for constant-curvature $T^{2}$-bundles over $T^{2}$).
  2. 2.
    The metric $g_{M}$ and connection forms $\Theta^{(\pm)}$ are globally defined on $M$.
  3. 3.
    The orthogonal splitting $T_{p}M=\mathcal{H}_{p}\oplus\mathcal{V}_{p}$ induced by $g_{M}$ is global.
These conditions hold regardless of the triviality of the underlying principal bundle, as long as the curvatures are constant (Assumption 2.1). The vanishing total Chern number $c^{\rm tot}_{1}=0$ ensures the existence of a global diagonal section, which is sufficient for the physical interpretation but not required for the mathematical proofs.
@@ 2.1 Kaluza–Klein metric on the extended space
Assumption 2.1 (Constant Berry curvatures).
In the following, we assume that the Berry curvature 2-forms are constant multiples of the Brillouin-zone volume form:
$F^{(\pm)}=c^{(\pm)}\,\mathrm{vol}_{\mathrm{BZ}},\qquad c^{(\pm)}\in\mathbb{R}\setminus\{0\},$
where $\mathrm{vol}_{\mathrm{BZ}}=dk_{x}\wedge dk_{y}$. The constants $c^{(\pm)}$ are related to the Chern numbers by $c_{1}^{(\pm)}=\frac{1}{2\pi}\int_{T^{2}}F^{(\pm)}=\frac{\operatorname{Area}(T^{2})}{2\pi}\,c^{(\pm)}$. This condition is satisfied in a wide range of experimentally relevant SOC BEC configurations, including linear Rashba–Dresselhaus couplings and uniform synthetic magnetic fields.
@@ 2.2 Metric connection with prescribed torsion
Define the three-form
$T^{b}:=F^{(+)}\wedge\Theta^{(+)}+F^{(-)}\wedge\Theta^{(-)}\in\Omega^{3}(M),$
where $\Theta^{(\pm)}=d\phi_{\pm}+\pi^{*}A^{(\pm)}$ are the horizontal 1-forms defined in (2.1). Note that $F^{(\pm)}\wedge A^{(\pm)}=0$ for dimensional reasons, since both forms are pullbacks from the 2-dimensional base $T^{2}_{\mathrm{BZ}}$. $T^{b}$ is the trace $\operatorname{Tr}(\mathcal{F}\wedge\boldsymbol{\Theta})$ with respect to the Killing form on $\mathfrak{u}(1)_{+}\oplus\mathfrak{u}(1)_{-}$, where $\boldsymbol{\Theta}=\Theta^{(+)}\otimes t_{+}+\Theta^{(-)}\otimes t_{-}$ is the total connection form on $M$ and $\mathcal{F}=F^{(+)}\otimes t_{+}+F^{(-)}\otimes t_{-}$.
@@ 2.3 Parallel forms in the Kaluza–Klein geometry
Lemma 2.5 (Triviality of the obstruction kernel $\mathcal{K}$ under constant curvature).
Let $g_{M}$ be the Kaluza–Klein metric (2.1) on $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, and let Assumption 2.1 hold. Then the obstruction kernel $\mathcal{K}$ vanishes identically:
$\mathcal{K}\;:=\;\bigl(\mathcal{V}_{2,1}\cap(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}))\bigr)\oplus\bigl(\mathcal{V}_{1,2}\cap(\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{2}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}))\bigr)\;=\;\{0\},$
where $\mathcal{V}_{2,1}$ and $\mathcal{V}_{1,2}$ are the spans of the harmonic representatives of the mixed Künneth components of the torsion class $\omega]$.
@@
Under Assumption 2.1, the cohomology classes of the Berry curvatures satisfy $F^{(-)}]=\lambda[F^{(+)}]$ with $\lambda=c^{(-)}/c^{(+)}$. Consequently, the mixed part of the torsion class is represented by the simple tensor
$\xi_{0}=\operatorname{vol}_{\mathrm{BZ}}\otimes\bigl(d\phi_{+}+\lambda d\phi_{-}\bigr)\in\mathcal{V}_{2,1}.$
Thus $\mathcal{V}_{2,1}=\operatorname{span}\{\xi_{0}\}$.
The space of parallel $1$-forms on the fibre, $\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$, consists of those linear combinations $\eta=c_{+}d\phi_{+}+c_{-}d\phi_{-}$ for which the corresponding vertical $1$-form $c_{+}\Theta^{(+)}+c_{-}\Theta^{(-)}$ is $g_{M}$-parallel. A direct computation (provided in Appendix A) shows that such an $\eta$ is parallel if and only if the constants satisfy the linear relation
$c_{+}+\lambda c_{-}=0.$
Now, an element of $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$ has the form
$\operatorname{vol}_{\mathrm{BZ}}\otimes(c_{+}d\phi_{+}+c_{-}d\phi_{-})\qquad\text{with}\qquad c_{+}+\lambda c_{-}=0.$
For this element to belong to $\mathcal{V}_{2,1}$, it must be proportional to $\xi_{0}$; hence there exists $t\in\mathbb{R}$ such that
$c_{+}=t,\qquad c_{-}=t\lambda.$
Substituting into condition (*) gives
$t+\lambda(t\lambda)=t(1+\lambda^{2})=0,$
which, for real $\lambda$, forces $t=0$ and therefore $c_{+}=c_{-}=0$. Hence the intersection $\mathcal{V}_{2,1}\cap(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}))$ is trivial. An analogous argument shows that the $(1,2)$-component intersection is also trivial, completing the proof. ∎
Remark 2.6 (Physical relevance of constant Berry curvatures).
Assumption 2.1 covers a wide range of experimentally relevant SOC BEC configurations. In Raman-induced spin–orbit coupling, the Berry curvature often becomes constant (or piecewise constant) in large regions of the Brillouin zone. Moreover, even when the curvature is not strictly constant, the topological lower bound depends only on the cohomology class $F^{(\pm)}]$, not on its pointwise values. One can always deform the connection within its cohomology class to a connection with constant curvature without changing the topological data that enter the bound. Therefore, the results derived under this assumption capture the essential topological obstruction present in generic SOC BECs.
@@ 3 Cohomology of the Extended Space and the Mixed Tensor Rank
In the following analysis, we work under Assumption 2.1 (constant Berry curvatures). With the geometric structure $(M,g_{M},\nabla^{C})$ established in Section 2, we now analyze the cohomological data that enter the general lower bound. Recall that the torsion three-form $T^{\flat}\in\Omega^{3}(M)$ is not necessarily closed, but its de Rham cohomology class $T^{\flat}]\in H^{3}(M;\mathbb{R})$ is well defined. Let $\omega_{h}$ be the unique $g_{M}$-harmonic representative of this class; then
@@ 3.1 Künneth decomposition of $H^{3}(M)$
Lemma 3.1 (Explicit harmonic representative).
Under Assumption 2.1, the unique $g_{M}$-harmonic representative of the torsion class $\omega]$ is given by
$\omega_{h}=\operatorname{vol}_{\mathrm{BZ}}\wedge\bigl(a_{+}\Theta^{(+)}+a_{-}\Theta^{(-)}\bigr),$
where $a_{\pm}$ are real constants satisfying $a_{-}=\lambda a_{+}$ with $\lambda=c^{(-)}/c^{(+)}$.
Proof.
Let $\eta=\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}\Theta^{(+)}+a_{-}\Theta^{(-)})$ with $a_{-}=\lambda a_{+}$.
Closedness:
$d\eta=\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}d\Theta^{(+)}+a_{-}d\Theta^{(-)})=\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}F^{(+)}+a_{-}F^{(-)})=(a_{+}c^{(+)}+a_{-}c^{(-)})\operatorname{vol}_{\mathrm{BZ}}\wedge\operatorname{vol}_{\mathrm{BZ}}=0.$
Coclosedness: In the orthonormal coframe $(dx,dy,\Theta^{(+)},\Theta^{(-)})$,
$\eta=a_{+}\Theta^{(-)}-a_{-}\Theta^{(+)},\quad d(*\eta)=a_{+}F^{(-)}-a_{-}F^{(+)}=(a_{+}c^{(-)}-a_{-}c^{(+)})\operatorname{vol}_{\mathrm{BZ}}.$
Since $a_{-}=\lambda a_{+}$ and $c^{(-)}=\lambda c^{(+)}$, we have $a_{+}c^{(-)}-a_{-}c^{(+)}=0$, hence $\delta\eta=*d(*\eta)=0$.
Thus $\eta$ is harmonic. Uniqueness follows from the Hodge theorem. ∎
@@
The representatives $d\phi_{\pm}$ are harmonic with respect to the standard metric on the circles, but not with respect to $g_{M}$. The $g_{M}$-harmonic representatives are the horizontal forms $\Theta^{(\pm)}$. Since $d\phi_{\pm}]=[\Theta^{(\pm)}]$ in $H^{1}(S^{1}_{\phi_{\pm}})$, the cohomological decomposition (3.1) remains valid when we replace $d\phi_{\pm}$ by $\Theta^{(\pm)}$.
@@ 3.2 Mixed tensor rank of the torsion class
For the class $\omega]$ coming from the physical torsion, the two-forms $\alpha_{+},\alpha_{-}$ in (3.1) are precisely the (cohomology classes of the) Berry curvature two-forms $F^{(+)},F^{(-)}$ of the two independent $U(1)$ sectors. A key observation is that $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$. Therefore, any two non-zero classes in this space are necessarily cohomologically proportional. Under Assumption 2.1, we have
$F^{(-)}]=\lambda[F^{(+)}]\quad\text{in }H^{2}(T^{2}_{\mathrm{BZ}}),\qquad\lambda=c^{(-)}/c^{(+)}.$
@@ 3.3 Vanishing of the obstruction kernel $\mathcal{K}$
In our geometry, we have $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\,\mathrm{vol}_{\mathrm{BZ}}$ (the space of parallel two-forms on the flat torus). The obstruction kernel $\mathcal{K}$ is shown to vanish by Lemma 2.5 under Assumption 2.1. The proof of Lemma 2.5 demonstrates that no non-zero element of the form $\operatorname{vol}_{\mathrm{BZ}}\otimes(c_{+}d\phi_{+}+c_{-}d\phi_{-})$ with $c_{+}+\lambda c_{-}=0$ can belong to $\mathcal{V}_{2,1}=\operatorname{span}\{\operatorname{vol}_{\mathrm{BZ}}\otimes(d\phi_{+}+\lambda d\phi_{-})\}$.
Thus we have